OpenStudy (anonymous):

If (2^x)(2^-4x) = 1/8, then x = ?

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

x=\[2^{-3x}=2^{-3}, x=1\]

OpenStudy (anonymous):

did u got answer

OpenStudy (anonymous):

the left simplifies to 2^(-3x) and the right can be rewritten as 2^(-3) so its 2^(-3x)=2^(-3) so x=1

OpenStudy (anonymous):

yes it is but i not really understand. Is it 1/2^3 = 1/8 so x = 1?

OpenStudy (anonymous):

1/2^-3x=1/8 so x = 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok I understand now. Thanks